OpenStudy (anonymous):

Give the name and structure of products formed when but-2-ene (H3C-CH=CH-CH3) is in the presence of a nickel dihydride catalyst (H2/Ni) and give the name and structure of products formed when but-2-ene (H3C-CH=CH-CH3) is in the presence of acidified water (H3O+ or H2O/H=)

OpenStudy (anonymous):

1) H3C-CH=CH-CH3 + H2 (using nickel as a catalyst) = H3C-CH2-CH2-CH3 (butane) 2) H3C-CH=CH-CH3 + H2O (using acid as a catalyst) = H3C-CH-(OH)-CH2-CH3 (2-butanol)